# Multiband Gravitational Wave Detection Prospects for M31 UCXB-1 System in Low and Middle Frequency Band

### v1

@@ III.1 GW Signal
where $P(n)$ ($n=1,2,3,\cdots$) is the angle-averaged GW power radiated in the $n$th harmonic from the eccentric binary given by (1963PhRv..131..435P; 2015PhRvD..92f3010H; maggiore2008gravitational)
@@
and $J_{n}$ is the $n$th order Bessel function. And the RMS amplitude of the angle-averaged GW strain in the $n$th harmonic is (2015PhRvD..92f3010H)
@@ III.2 S/N
If adopting the matched-filtering method, the S/N $\varrho$ of this GW signal from eccentric binary can be roughly estimated as (2025ApJ...978..104G)
@@ IV Conclusions
GW, serving as a complementary probe to electromagnetic observations, provide information inaccessible to electromagnetic observation method alone, such as constraints on the primary mass $m_{1}$ for M31 UCXB-1 system. As a representative and vital multi-messenger source, M31 UCXB-1 can be observed not only through both GW and electromagnetic channels, but also via GW across different frequency bands (low and middle frequency band), thereby leveraging the strengths of detectors operating in distinct parts of the GW spectrum. Given $m_{1}=20M_{\odot}$, BBO can constrain the relative uncertainty of the $\Delta m_{1}/m_{1}$ parameter to $\sim 30\%(20\%)$ during $T_{\rm obs}=4(10)$ yr. Multiband joint observations of GW sources can help enhance the S/N of the GW signal, improve localization accuracy, and reduce parameter estimation errors (2020MNRAS.496..182L; 2023MNRAS.522.2951Z; 2025arXiv251213151Z). Multiband joint observations in low and middle frequency band represent a promising approach worthy of consideration in future studies of M31 UCXB-1 or similar short-period UCXB with $T_{\rm orb}<10^{3}$ s. In this work, we consider a joint network of all low- and middle-frequency GW detectors, referred to as the TLTDB network. During a 10(4)-year observation period, the powerful TLTDB detector network can achieve a detection S/N of 8 as long as the GW source with $m_{1}>5.4(10.6)\,M_{\odot}$.
@@ Appendix B Merger time & semimajor axis
If its orbit is eccentric, the timescale for an eccentric binary with semimajor axis $a$ and eccentricity $e$ decaying to merger by the GW radiation is (1964PhRv..136.1224P; maggiore2008gravitational),

### v2

@@ III.1 GW Signal
where $P(n)$ ($n=1,2,3,\cdots$) is the angle-averaged GW power radiated in the $n$th harmonic from the eccentric binary given by (Peters and Mathews, 1963; Huerta et al., 2015; Maggiore, 2008)
@@
and $J_{n}$ is the $n$th order Bessel function. And the RMS amplitude of the angle-averaged GW strain in the $n$th harmonic is (Huerta et al., 2015)
@@ III.2 S/N
If adopting the matched-filtering method, the S/N $\varrho$ of this GW signal from eccentric binary can be roughly estimated as (Guo et al., 2025)
@@ IV Conclusions
GW, serving as a complementary probe to electromagnetic observations, provide information inaccessible to electromagnetic observation method alone, such as constraints on the primary mass $m_{1}$ for M31 UCXB-1 system. As a representative and vital multi-messenger source, M31 UCXB-1 can be observed not only through both GW and electromagnetic channels, but also via GW across different frequency bands (low and middle frequency band), thereby leveraging the strengths of detectors operating in distinct parts of the GW spectrum. Given $m_{1}=20M_{\odot}$, BBO can constrain the relative uncertainty of the $\Delta m_{1}/m_{1}$ parameter to $\sim 30\%(20\%)$ during $T_{\rm obs}=4(10)$ yr. Multiband joint observations of GW sources can help enhance the S/N of the GW signal, improve localization accuracy, and reduce parameter estimation errors (Liu et al., 2020a; Zhao et al., 2023, 2025). Multiband joint observations in low and middle frequency band represent a promising approach worthy of consideration in future studies of M31 UCXB-1 or similar short-period UCXB with $T_{\rm orb}<10^{3}$ s. In this work, we consider a joint network of all low- and middle-frequency GW detectors, referred to as the TLTDB network. During a 10(4)-year observation period, the powerful TLTDB detector network can achieve a detection S/N of 8 as long as the GW source with $m_{1}>5.4(10.6)\,M_{\odot}$.
@@ Appendix B Merger time & semimajor axis
If its orbit is eccentric, the timescale for an eccentric binary with semimajor axis $a$ and eccentricity $e$ decaying to merger by the GW radiation is (Peters, 1964; Maggiore, 2008),